OpenStudy (anonymous):

Evaluate the limit as x approaches 0 for {3-(9-x)^.5}/x

OpenStudy (anonymous):

Factor an x from the top and bottom and cancel them out.

OpenStudy (anonymous):

use LHopital's rule

OpenStudy (anonymous):

Oh, sorry I misread. Can't get that x out of the parens

OpenStudy (anonymous):

differentiate numerator and denominator separately

OpenStudy (anonymous):

that x on the top is inside of a radical

OpenStudy (anonymous):

1/(2sqrt(9-x) will be the result. then plug x = 0 1/(2sqrt(9) = 1/6

OpenStudy (anonymous):

HOLD ON Kdabr what calculus are you in.... helping tutors, he can't use l'hopitals unless he's in calc 2

OpenStudy (anonymous):

or unles he's been taught it

OpenStudy (anonymous):

Im in precal haha, about to enter calc BC

OpenStudy (anonymous):

So probably not l'hopitals since he's just getting into limits

OpenStudy (anonymous):

helping tutors got the right answer but I have no idea what he did

OpenStudy (anonymous):

here kdabr... you're goingto have to rationalize it.. and yes kdabr, he used l'hopitals rule which is taught after calculus 1 as there is more rules with it that can make it tricky..I have a problem much ilke it in my book i'll send you it now

OpenStudy (anonymous):

I have use the L'hopital's rule.. and since you are in precalc... rationalisation could work better for your grade

OpenStudy (anonymous):

that page looks spot on

OpenStudy (anonymous):

your problem is very similar to this problem if you were to mimic what it looks like in my book you don't necessarily have to tho

OpenStudy (anonymous):

I tried doing that already and ended up with x/(3x+sqrt(9-x))

OpenStudy (anonymous):

now if you want to dabble into l'hopitals to perhaps check yourself i could send you that also but if you were to solve usinng L'H you'd probably get 0 pts

OpenStudy (anonymous):

you are correct about that outkast

OpenStudy (anonymous):

and also it does not work in all cases only certain indeterminate forms

OpenStudy (anonymous):

i'm hoping polpak is doing the problem

OpenStudy (anonymous):

as i amnot lol

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{3-\sqrt{9-x}}{x} \]\[= \lim_{x\rightarrow 0}\frac{3-\sqrt{9-x}}{x} \cdot \frac{3+\sqrt{9-x}}{3+\sqrt{9-x}}\]\[=\lim_{x\rightarrow0}\frac{9-(9-x)}{x(3+\sqrt{9-x})}\]\[=\lim_{x\rightarrow0} \frac{9-9 + x}{x(3+\sqrt{9-x})}\]\[=\lim_{x\rightarrow0} \frac{x}{x(3+\sqrt{9-x})}\]\[=\lim_{x\rightarrow0} \frac{1}{3+\sqrt{9-x}}\]\[= \frac{1}{3 + \sqrt{9 + 0}} = \frac{1}{3+3} = \frac{1}{6}\]

OpenStudy (anonymous):

yes indeed what he did you didn't pull out an x and cancel the top... it should all cancel

OpenStudy (anonymous):

You forgot your factor of x on the bottom

OpenStudy (anonymous):

but why do those x's become ones?

OpenStudy (anonymous):

to cancel with the x on the top.

OpenStudy (anonymous):

oh of course

OpenStudy (anonymous):

Because x/x = 1 as long as x isn't 0 and we're approaching 0, we don't actually care what happens when we get there.

OpenStudy (anonymous):

did you want l'hopitals for a check or idk even if i'td make sense to you as you probably dont know how to differentiate?

OpenStudy (anonymous):

I think he has the answer in a key. Just didn't know how to find it.